OpenStudy (anonymous):

Whats (-4,6);m=3/4 in point slope form) can you help me remember the steps to this

OpenStudy (opcode):

\[y - y_1 = m(x - x_1)\] \[m = slope\]

OpenStudy (opcode):

You need two points, this only provides one. \(O.o?\)

OpenStudy (anonymous):

thats all the test says is write this equation in point slope form (-4,6); m=3/4

OpenStudy (opcode):

Is the question multiple choice?

OpenStudy (anonymous):

yes but i cant ask for answers on here

OpenStudy (anonymous):

I believe the eqution would be y-6=(3/4)(x--4)

zepdrix (zepdrix):

You're given the point \(\Large\bf\sf (x_1,y_1)\) and the slope, \(\Large\bf\sf \dfrac{3}{4}\). Just plug them into the formula that OP posted. Yah like chromer did :)

OpenStudy (anonymous):

ok thanks every1